Find the period and amplitude of y = -5 - sin 4x.
Answer #1

His Solution y=-5-sinux y=-sinux-S y = Asin (Bx+()+D A = -1 B = 4 c=0 D=-5 1 Amplitude IA 1-1 = 1 period= an B 79 Ft 2T perio